Christy is 3 years older than Jamie.  The sum of their ages is 23.  How old is each?
Found 2 solutions by Ave, stanbon:
Answer by Ave(106)

Jamie = x
Christy = x+3
(x)+(x+3)= 23
2x+3=23
2x=20
x=10
Jamie =10
Christy =10+3=13

Answer by stanbon(75887)

Christy is 3 years older than Jamie. The sum of their ages is 23. How old is each?
-----
Equations:
c = j + 3
c + j = 23
---------
Substitute for "c" and solve for "j":
j+3 + j = 23
2j = 20
j = 10 yrs old (Jamie's age now)
-------
Solve for "c":
c = j + 3
c = 13 yrs old (Christy's age now)
===================
Cheers,
Stan H.
===================